# GPa Pressure Imaging Using Nanodiamond Quantum Sensors

### v1

@@ I Introduction
Pressure is a fundamental thermodynamic variable and has long been a key tuning parameter in condensed matter physics [1]. Its application is of particular interest, as it can induce phase transitions and lead to emergent electronic properties. A notable example is oxygen, which becomes metallic and superconducting under high pressure [2, 3], despite being a gas under ambient conditions. More recently, hydrogen-rich materials and nickelates have attracted significant attention as high-temperature superconductors under pressure [4, 5, 6, 7, 8].
Progress in high-pressure research has been driven largely by advances in experimental techniques. The diamond anvil cell (DAC) is now an essential tool, capable of generating pressures up to hundreds of gigapascals while allowing various in situ measurements [9]. Accurate pressure determination within the DAC sample chamber is crucial in such experiments. Common methods include ruby fluorescence [10, 11, 12], diamond Raman spectroscopy [13, 14, 15], and diffraction-based techniques such as X-ray [16] and neutron diffraction [17, 18, 19]. Each approach offers unique advantages depending on the experimental context.
Quantum sensing based on nitrogen-vacancy (NV) centers in diamonds has recently emerged as a powerful technique for local measurement of physical quantities [20, 21, 22, 23, 24, 25]. Their electron spin resonance can be optically detected via photoluminescence (PL) under simultaneous laser and microwave irradiation, a technique known as optically detected magnetic resonance (ODMR).
This enables non-invasive measurement of magnetic fields, temperature, and pressure. Embedding NV centers near the culet surface of a DAC facilitates probing of pressure and magnetic fields in close proximity to the sample, which has been difficult using conventional techniques. This approach has recently been employed to study novel phenomena such as magnetic phase transitions in iron and superconducting transitions in hydrides [26, 27, 28, 29, 30, 31, 32].
In this context, several studies have proposed an alternative approach in which microdiamonds [32, 33] or nanodiamonds (NDs) [30, 34] containing NV centers are embedded directly into the DAC sample chamber. Unlike the method of embedding NV centers into the culet, this method places NV centers within the chamber, allowing more direct probing of the sample’s local environment. As hard diamond particles are introduced into the chamber with pressure-transmitting medium (PTM), it is crucial to identify embedding strategies compatible with the desired pressure conditions. Moreover, the use of ND ensembles introduces crystal orientation variations that complicate ODMR spectra, highlighting the need for analytical techniques suitable for optical imaging with CMOS cameras.
@@ II.1 NV Centers
An NV center consists of a substitutional nitrogen atom adjacent to a vacancy, as illustrated in Fig. 1(a). This defect has $C_{3v}$ symmetry, with its axis, known as the NV axis, aligned along one of the $\langle 111\rangle$ crystallographic directions. The negatively charged NV center has an electron spin of $S=1$ [35], and we refer to this charge state simply as the NV center in the following. The spin energy levels in the electronic ground state are shown in Fig. 1(b). The spin states with $m_{S}=0$, $-1$, and $+1$ correspond to projections of spin angular momentum along the NV axis. In the absence of a magnetic field and at ambient pressure, the $m_{S}=0$ and $m_{S}=\pm 1$ states are split by the zero-field splitting $D_{\mathrm{gs}}=2.87~{}\mathrm{GHz}$[23]
@@
The spin states of the NV center can be polarized or initialized to the $m_{S}=0$ state via optical excitation and read out through PL, owing to spin-dependent optical transitions. This enables ODMR spectroscopy, where electron spin resonance is measured by monitoring PL intensity under continuous laser illumination while sweeping the microwave frequency. Figure 1(c) shows simulated ODMR spectra. The vertical axis indicates the PL ratio, which is defined as the PL intensity with microwave irradiation divided by that without. A dip appears when the microwave frequency matches the spin resonance. At zero magnetic field and ambient pressure, a dip appears at the zero-field splitting $D_{\mathrm{gs}}=2.87~{}\mathrm{GHz}$. Shifts in energy levels lead to changes in the shape and position of the ODMR spectrum. Thus, physical quantities such as pressure, magnetic field, and temperature can be extracted from ODMR measurements [21, 27, 36, 37, 38, 39]. This quantum sensing technique, based on optical detection, is compatible with wide-field microscopy for imaging applications. [40]
@@ II.2 Pressure Sensing via ODMR Spectra
In this subsection, we describe the stress dependence of an individual NV center in a ND. The ODMR spectra of NV centers in ND ensembles are discussed in the following subsection. To model the stress dependence, we consider the effective spin Hamiltonian [26, 27, 38, 41, 42, 43, 44, 45]. In this study, we neglect external perturbations such as magnetic fields and employ the following spin Hamiltonian:
@@
where $\sigma_{ij}$ are components of the local stress tensor. The coefficients are known as $\alpha_{1}=8.6$ MHz/GPa, $\alpha_{2}=-1.95$ MHz/GPa, $\beta_{1}=-2.5$ MHz/GPa, and $\beta_{2}=-4.5$ MHz/GPa [27, 30]. The resonance frequencies $f_{\pm}$, obtained by diagonalizing the Hamiltonian, are
@@ II.3 ODMR Spectrum under Uniaxial Stress
The stress environment inside the DAC sample chamber is often not purely hydrostatic. Instead, as illustrated in Fig. 2(a), a uniaxial stress condition frequently develops, characterized by a higher pressure along the loading axis ($P_{Z}$) than in the transverse ($P_{\perp}$) directions [27, 30, 31, 34, 46]. In the laboratory coordinate system $(X,Y,Z)$, where the DAC loading axis is defined as the $Z$-axis, such uniaxial stress environment is expressed as follows:
@@
To estimate the ODMR spectrum of ND ensemble measurements, the signals from all possible NV center orientations $(\theta,\phi,\psi)$ must be summed. This summation must account for both variations in the ODMR spectra and differences in PL intensity, arising from the anisotropy of correction efficiency $A$ and absorption efficiency $\kappa$[47, 48].
@@
×$\times M(\theta,\phi)\,\kappa(\theta,\phi)\,A(\theta,\theta_{\text{max}})]+\delta,$ (13)
where $I_{\mathrm{NDs}}$ is the ODMR spectrum of the ND ensemble measurement, $I_{\mathrm{NV}}$ is the ODMR spectrum shape from the NV center at a specific orientation $(\theta,\phi,\psi)$, $D$ denotes the integration range determined by the numerical aperture (NA) of the objective lens, where $\mathrm{NA}/n_{d}=\sin(\theta_{\mathrm{max}})$ with $n_{d}$ the refractive index of diamond[49], and $\delta$ denotes an offset term.
@@
$A(\theta,\theta_{\text{max}})$ $\propto\frac{\pi}{12}\left[32-\cos\theta_{\text{max}}\left(31+\cos(2\theta_{\text{max}})\right.\right.$
@@
To account for not only the optical measurement anisotropy but also the directional dependence of the microwave magnetic field, we introduce a weighting factor $M$. The ODMR contrast of an NV center increases with the amplitude of the microwave magnetic field component perpendicular to the NV axis. The factor $M$ is based on the assumption that, when the microwave amplitude is weak, the contrast depends quadratically on the microwave amplitude. [30, 50, 51]
@@
Under the assumption of a uniaxial stress environment, the fitting parameters are only five: $C$, $d\nu$, $P_{Z}$, $P_{\perp}$, and $\delta$. The small number of fitting parameters makes it suitable for imaging measurements that require fitting a large number of ODMR spectra. Under hydrostatic conditions, $P=P_{Z}=P_{\perp}$, which reduces the number of fitting parameters and simplifies the spectral shape to the form given by Eq. (6).
A similar model has been proposed by Shelton et al. [30], but our approach subsumes both hyperfine and electric field broadening into a single Lorentzian linewidth $d\nu$, while also incorporating the anisotropy of correction efficiency $A(\theta,\theta_{\max})$ and absorption efficiency $\kappa(\theta,\phi)$ [47, 48]. This description retains sufficient information for ODMR spectrum analysis and makes our model particularly well suited for rapid, large-area pressure imaging measurements.
@@ III.1 Experimental setup
The pressure is applied by sandwiching the sample between the upper and bottom diamond anvils [Fig. 3(c)]. We use a pair of diamond anvils with a culet diameter of $260\,\mathrm{\mu m}$ and a rhenium (Re) gasket with an initial thickness of $250\,\mathrm{\mu m}$. It is pre-indented to a thickness of $42\,\mathrm{\mu m}$ using the diamond anvils, and a $100\,\mathrm{\mu m}$ diameter hole is drilled at the center using an infrared laser to form the sample chamber. The DAC sample chamber is filled with sodium chloride ($\mathrm{NaCl}$) as a PTM, along with 50 nm- sized NDs (Adámas Nanotechnologies, NDNV50nmHi10ml) for measurement and a ruby crystal of about $10~{}\mu$m in size for pressure calibration. A small iron sample (4N purity, Nilaco Co.; $\sim 10\,\mu\mathrm{m}^{t}$) is also embedded for future magnetic field evaluation.
@@ III.2.1 Single-layer PTM
Figure 3(d) shows a schematic cross-sectional view of the sample chamber. Figure 3(f) is a microscope image taken from the top of the chamber. The gray dashed line indicates the boundary of the sample chamber, and the pink circle marks the location of the ruby used for pressure calibration. The pressure obtained from the ruby fluorescence is $20.6~{}\mathrm{GPa}$ [11]. In this configuration, the NDs, the ruby, and iron sample are in contact with the upper diamond anvil.
@@ III.2.2 Double-layer PTM
Note that both the NDs and NaCl are too small and transparent to be visible in the optical microscope images. The pressure obtained from the ruby fluorescence is $20.0~{}\mathrm{GPa}$ [11]. In this configuration, an additional NaCl layer is placed above the ND layer and measurement samples, so that the NDs, ruby, and iron are not in contact with the culet surface of the upper diamond anvil.
@@ IV.1 Single-layer PTM Configuration
Although accurate pressure determination by ruby fluorescence is technically difficult under non-hydrostatic conditions, [52, 53, 54] the pressure estimated from NV centers agrees, at least in order of magnitude, with the ruby result, supporting the validity of the present model.
@@ Acknowledgements
Figure S1: Results from the single-layer pressure medium experiment. We analyzed this result using both the hydrostatic model and the uniaxial stress model. (a,c) Spatial distribution of pressure [(a)uniaxial stress model:$(P_{Z}+2P_{\perp})/3$ (c)hydrostatic model:$P$] inside the sample chamber. The white outline indicates the sample chamber, the black box marks the iron sample, and the red circle marks the ruby. Black pixels represent regions with insufficient signal intensity. (b,d) Histogram of pressure [(b)uniaxial stress model:$(P_{Z}+2P_{\perp})/3$ (d)hydrostatic model:$P$] within the sample chamber. The red solid line indicates the pressure value obtained from ruby fluorescence ($20.6\,\mathrm{GPa}$). (e)Residuals between the experimental ODMR contrast and the two models: hydrostatic (green line) and uniaxial stress (blue line). The red box plots indicate the experimental uncertainty in the ODMR contrast.

### v2

@@ I Introduction
Pressure is a fundamental thermodynamic variable and has long been a key tuning parameter in condensed matter physics [Mao2018-sv]. Its application is of particular interest, as it can induce phase transitions and lead to emergent electronic properties. A notable example is oxygen, which becomes metallic and superconducting under high pressure [ShimizuNature1998, ShimizuJPSJ2005], despite being a gas under ambient conditions. More recently, hydrogen-rich materials and nickelates have attracted significant attention as high-temperature superconductors under pressure [Drozdov2015-pm, Einaga2016-mv, DrozdovNature2019, SomayazuluPRL2019, SunNature2023].
Progress in high-pressure research has been driven largely by advances in experimental techniques. The diamond anvil cell (DAC) is now an essential tool, capable of generating pressures up to hundreds of gigapascals while allowing various in situ measurements [JayaramanRMP1983]. Accurate pressure determination within the DAC sample chamber is crucial in such experiments. Common methods include ruby fluorescence [Mao1976-kp, Mao1978-zu, Mao1986-kc], diamond Raman spectroscopy [Hanfland1985-pz, Eremets2003-tq, Akahama2004-gj], and diffraction-based techniques such as X-ray [Zha2000-ez] and neutron diffraction [Jorgensen1984-ik, Jorgensen1985-nq, Besson1992-xn]. Each approach offers unique advantages depending on the experimental context.
Quantum sensing based on nitrogen-vacancy (NV) centers in diamonds has recently emerged as a powerful technique for local measurement of physical quantities [Taylor2008, Balasubramanian2008, Rondin2014, Schirhagl2014-dw, DegenRMP2017, Casola2018]. Their electron spin resonance can be optically detected via photoluminescence (PL) under simultaneous laser and microwave irradiation, a technique known as optically detected magnetic resonance (ODMR).
This enables non-invasive measurement of magnetic fields, temperature, and pressure. Embedding NV centers near the culet surface of a DAC facilitates probing of pressure and magnetic fields in close proximity to the sample, which has been difficult using conventional techniques. This approach has recently been employed to study novel phenomena such as magnetic phase transitions in iron and superconducting transitions in hydrides [Steele2017-po, Hsieh2019-op, Lesik2019-vw, Ho2020-cb, Shelton2024-kl, Wang2024-rf, Yip2019-of].
In this context, several studies have proposed an alternative approach in which microdiamonds [Yip2019-of, Shang2019-ol] or nanodiamonds (NDs) [Shelton2024-kl, Ho2023-dp] containing NV centers are embedded directly into the DAC sample chamber. Unlike the method of embedding NV centers into the culet, this method places NV centers within the chamber, allowing more direct probing of the sample’s local environment. As hard diamond particles are introduced into the chamber with pressure-transmitting medium (PTM), it is crucial to identify embedding strategies compatible with the desired pressure conditions. Moreover, the use of ND ensembles introduces crystal orientation variations that complicate ODMR spectra, highlighting the need for analytical techniques suitable for optical imaging with CMOS cameras.
@@ II.1 NV Centers
An NV center consists of a substitutional nitrogen atom adjacent to a vacancy, as illustrated in Fig. 1(a). This defect has $C_{3v}$ symmetry, with its axis, known as the NV axis, aligned along one of the $\langle 111\rangle$ crystallographic directions. The negatively charged NV center has an electron spin of $S=1$ [Loubser1978-lc], and we refer to this charge state simply as the NV center in the following. The spin energy levels in the electronic ground state are shown in Fig. 1(b). The spin states with $m_{S}=0$, 1-1, and +1+1 correspond to projections of spin angular momentum along the NV axis. In the absence of a magnetic field and at ambient pressure, the $m_{S}=0$ and $m_{S}=\pm 1$ states are split by the zero-field splitting $D_{\mathrm{gs}}=2.87~\mathrm{GHz}$[Schirhagl2014-dw]
@@
The spin states of the NV center can be polarized or initialized to the $m_{S}=0$ state via optical excitation and read out through PL, owing to spin-dependent optical transitions. This enables ODMR spectroscopy, where electron spin resonance is measured by monitoring PL intensity under continuous laser illumination while sweeping the microwave frequency. Figure 1(c) shows simulated ODMR spectra. The vertical axis indicates the PL ratio, which is defined as the PL intensity with microwave irradiation divided by that without. A dip appears when the microwave frequency matches the spin resonance. At zero magnetic field and ambient pressure, a dip appears at the zero-field splitting $D_{\mathrm{gs}}=2.87~\mathrm{GHz}$. Shifts in energy levels lead to changes in the shape and position of the ODMR spectrum. Thus, physical quantities such as pressure, magnetic field, and temperature can be extracted from ODMR measurements [Balasubramanian2008, Hsieh2019-op, Maze2008-or, Acosta2010-lt, Barfuss2019-db, Yamamoto2025-df]. This quantum sensing technique, based on optical detection, is compatible with wide-field microscopy for imaging applications. [Pham2011-bk]
@@ II.2 Pressure Sensing via ODMR Spectra
In this subsection, we describe the stress dependence of an individual NV center in a ND. The ODMR spectra of NV centers in ND ensembles are discussed in the following subsection. To model the stress dependence, we consider the effective spin Hamiltonian [Steele2017-po, Hsieh2019-op, Barfuss2019-db, Doherty2014-gg, Doherty2012-mi, Barson2017-zy, Udvarhelyi2018-if, felton2009hyperfine]. In this study, we neglect external perturbations such as magnetic fields and employ the following spin Hamiltonian:
@@
where $\sigma_{ij}$ are components of the local stress tensor. The coefficients are known as $\alpha_{1}=8.6$ MHz/GPa, $\alpha_{2}=-1.95$ MHz/GPa, $\beta_{1}=-2.5$ MHz/GPa, and $\beta_{2}=-4.5$ MHz/GPa [Hsieh2019-op, Shelton2024-kl]. The resonance frequencies $f_{\pm}$, obtained by diagonalizing the Hamiltonian, are
@@ II.3 ODMR Spectrum under Uniaxial Stress
The stress environment inside the DAC sample chamber is often not purely hydrostatic. Instead, as illustrated in Fig. 2(a), a uniaxial stress condition frequently develops, characterized by a higher pressure along the loading axis ($P_{Z}$) than in the transverse ($P_{\perp}$) directions [Hsieh2019-op, Shelton2024-kl, Wang2024-rf, Ho2023-dp, Bhattacharyya2024-oa]. In the laboratory coordinate system $(X,Y,Z)$, where the DAC loading axis is defined as the $Z$-axis, such uniaxial stress environment is expressed as follows:
@@
To estimate the ODMR spectrum of ND ensemble measurements, the signals from all possible NV center orientations $(\theta,\phi,\psi)$ must be summed. This summation must account for both variations in the ODMR spectra and differences in PL intensity, arising from the anisotropy of collection efficiency $A$ and absorption efficiency $\kappa$[Horowitz2012-nw, Tsukamoto2022-sb].
@@
$\times M(\theta,\phi)\,\kappa(\theta,\phi)\,A(\theta,\theta_{\text{max}})]+\delta,$ (13)
where $I_{\mathrm{NDs}}$ is the ODMR spectrum of the ND ensemble measurement, $I_{\mathrm{NV}}$ is the ODMR spectrum shape from the NV center at a specific orientation $(\theta,\phi,\psi)$, $D$ denotes the integration range determined by the numerical aperture (NA) of the objective lens, where $\mathrm{NA}/n_{d}=\sin(\theta_{\mathrm{max}})$ with $n_{d}$ the refractive index of diamond[thomas1993optical], and $\delta$ denotes an offset term.
@@
[…]
@@
To account for not only the optical measurement anisotropy but also the directional dependence of the microwave magnetic field, we introduce a weighting factor $M$. The ODMR contrast of an NV center increases with the amplitude of the microwave magnetic field component perpendicular to the NV axis. The factor $M$ is based on the assumption that, when the microwave amplitude is weak, the contrast depends quadratically on the microwave amplitude. [Shelton2024-kl, Dreau2011-gc, Jensen2013-nb]
@@
Under the assumption of a uniaxial stress environment, the fitting parameters are only five: $C$, $d\nu$, $_{Z}, PP_{\perp}$, and $\delta$. The small number of fitting parameters makes it suitable for imaging measurements that require fitting a large number of ODMR spectra. Under hydrostatic conditions, $P=P_{Z}=P_{\perp}$, which reduces the number of fitting parameters and simplifies the spectral shape to the form given by Eq. (6).
A similar model has been proposed by Shelton et al. [Shelton2024-kl], but our approach subsumes both hyperfine and electric field broadening into a single Lorentzian linewidth $d\nu$, while also incorporating the anisotropy of collection efficiency $A(\theta,\theta_{\max})$ and absorption efficiency $\kappa(\theta,\phi)$ [Horowitz2012-nw, Tsukamoto2022-sb]. This description retains sufficient information for ODMR spectrum analysis and makes our model particularly well suited for rapid, large-area pressure imaging measurements.
@@ III.1 Experimental setup
[…]
@@ III.2.1 Single-layer PTM
Figure 3(d) shows a schematic cross-sectional view of the sample chamber. Figure 3(f) is a microscope image taken from the top of the chamber. The gray dashed line indicates the boundary of the sample chamber, and the pink circle marks the location of the ruby used for pressure calibration. The pressure obtained from the ruby fluorescence is $20.6~\mathrm{GPa}$ [Mao1978-zu]. In this configuration, the NDs, the ruby, and iron sample are in contact with the upper diamond anvil.
@@ III.2.2 Double-layer PTM
Note that both the NDs and NaCl are too small and transparent to be visible in the optical microscope images. The pressure obtained from the ruby fluorescence is $20.0~\mathrm{GPa}$ [Mao1978-zu]. In this configuration, an additional NaCl layer is placed above the ND layer and measurement samples, so that the NDs, ruby, and iron are not in contact with the culet surface of the upper diamond anvil.
@@ IV.1 Single-layer PTM Configuration
Although accurate pressure determination by ruby fluorescence is technically difficult under non-hydrostatic conditions, [Piermarini1973-jb, Fujishiro1988-lc, He1995-nt] the pressure estimated from NV centers agrees, at least in order of magnitude, with the ruby result, supporting the validity of the present model.
@@ Acknowledgements
[…]